Cull Corporation uses a job-order costing system with a single plantwide predetermined overhead rate based on machine-hours. The company based its predetermined overhead rate for the current year on total fixed manufacturing overhead cost of $462,000, variable manufacturing overhead of $2.20 per machine-hour, and 60,000 machine-hours. The company has provided the following data concerning Job X455 which was recently completed:

Number of units in the job 20
Total machine-hours 80
Direct materials $ 940
Direct labor cost $ 2,240

The amount of overhead applied to Job X455 is closest to: (Round your intermediate calculations to 2 decimal places.)
Answer #1

Predetermine overhead rate = estimated overhead/estimated labor hour = (462000/60000)+2.2 = 9.90 per MH

The amount of overhead applied = Actual machine hour*Overhead rate = 80*9.9 = 792